DEC 17 OWL SHIFT SUMMARY

Rick Kessler (kessler@hep.uchicago.edu)
Sat, 18 Dec 1999 13:49:58 -0600

Shift Summary for eve shift, 17/Dec/99

On Shift: Pat T., Rick K.

Hours of beam available at all intensities: 7.3
Hours of beam available at nominal intensity: 7.0
Hours of E799 data (special runs not included):
Hours of special runs (laser scans, muon runs...): 7.0

Accounting of lost time: Starting/stopping runs.

========================================================================
Runs ended this shift:
Run 15238 - 2.8M events - BEAM_799
Run 15239 - 5.8M events - BEAM_799
Run 15240 - 2.5M events - BEAM_799
Run 15241 - in progress after NM2SR flip

Tapes brought to FCC this shift: ??? We know Sydney took some.

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E799 datataking

Plagued by incessant scaler alarm errors, which are due to the
SEM reading being off by one spill. So every time the intensity
changes, it matches the previous SEM with the current scalers,
and then goes nuts and makes noise.

MCR re-booted everything, and we re-started all atev1 and atev2
processes using the "sub_job.tcl" script ... but no success.
So I copied "scaler_check.tcl -> scaler_check.tcl_Dec17 and then
removed most of the scaler alarms.

BA3 hardlines were unplugged to reduce data vol readout time
from Cr 8.

Massive power outage around 11:30.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15:30 Shift starts with no beam.

15238 16:00 Start run.
16:10 Beam returns, but intensity starts off low (2E12).
We soon get many "Scaler out of range" alarms due
to scalers being too low. This is odd because we
expect the allowed range would be too low at low
intensity (see logbook 23 p.97). Eventually the
intensity reaches 9.5E12, but the alarms continue.
16:45 Sasha L. notices that the SEM reported by
scaler_screen is off by one spill. We page Ashkan.
17:15 On Ashkan's advice, we ask MCR to reboot their
SAD process. They do so.
17:33 Stop run 15238.

15239 17:39 Start run. SEM still off by 1 spill.
19:15 We finally give in and edit the scaler_check.map
file to turn off most of the alarms. Salah from
MCR shows up to investigate the spill mismatch.
Ch.13 reports the correct SEM, yet somehow sdaq
is getting it a spill late.
20:30 MCR guys leave to do secret stuff. Now the SEM
reported on the scaler screen agrees with Ch.13,
but the scaler rates in sdaq are still not right.
20:40 Stop run 15239. Beam gone for 15 min.

12240 20:45 Start run, no beam until spill 20. At spill 27
we stop triggers and restart all atev1 and atev2
processes. No good, still off by 1 spill.
22:30 Beam becomes intermittant, then goes away. Run stops.

23:15 STOP RUN to flip NM2SR [random carriage returns at
epciure window finally seemed to work]

12241 23:25 START new run, but no beam.
23:35 The power goess out, effectively ending our shift
========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

=> SEM reading is still one spill out of sync.
=> scaler alarms are disabled !

YO, RUN MANAGER: Don't forget to copy the nominal scaler_check file
back when the SEM-read gets fixed.

*** END of REPORT ***